OpenStudy (shylovesyou11):

Please help me with this math, i don't understand it =\ 1) What is the slope of the line that passes through the points (1, 3) and (5, –2)? 2) Select the equation of the line in slope-intercept form: 4x + 9y = 36

OpenStudy (anonymous):

-2-3=-5 over 5-1=4 so -5/4

OpenStudy (shylovesyou11):

Okay, Thank you! ^_^

OpenStudy (anonymous):

slpe intercept form =y=mx+b so y=-4/9x+4

OpenStudy (shylovesyou11):

Thank you sooo much!

OpenStudy (anonymous):

welcome :)